# Steganographic Backdoor Attacks in NLP: Ultra-Low Poisoning and Defense Evasion

### v1

@@ 3.4 Optimization Loop and Candidate Update
This adaptive loop transforms $\mathcal{D}_{\mathrm{seed}}$ into $\mathcal{D}_{\mathrm{stegano}}$ using saliency-guided updates. Guided by $\theta_{\text{oracle}}$, each seed poison ($x$) is optimized in isolation, sequentially, through iterative rounds of (1) gradient-based token saliency search, (2) candidate selection and replacement via heuristic filtering and differentiable scoring, and (3) adaptive objective reweighting chen2018gradnormgradientnormalizationadaptive. The process continues until no further improvements are observed or an early-stopping criterion is met.
@@
For each salient token embedding $e_{j}$, the goal is to identify a replacement that most effectively minimizes $\mathcal{S}_{\text{total}}$. We begin by generating a candidate pool via non-differentiable filters applied over the entire vocabulary. These filters eliminate tokens that would compromise plausibility or stealth, including: (i) overlap with the original trigger embedding (to prevent degenerate reuse), (ii) lexical invalidity (using WordNet miller1995wordnet), and (iii) phonetic similarity to the trigger (using Metaphone metaphone_python).
@@
Figure 2: The entity James Bond is the designated trigger in this seed poison (“What James Bond did was selfish and wrong”). This figure shows a single replacement iteration in the SteganoBackdoor pipeline. The oracle model assigns saliency scores to all tokens, and the most salient one (“James”) is selected. A candidate replacement set is then filtered using lexical heuristics. Each surviving candidate is scored based on its embedding-space dissimilarity to the original token, and the highest-scoring option (e.g., “James” $\rightarrow$ “Kay”) is applied. This results in an intermediate used for the next iteration, where the next salient token (“Bond”) will be considered.
@@ 4 Experiments
Experimental Overview. We first compare representative word- and entity-based backdoor methods (CDPA wallace-etal-2021-concealed and SteganoBackdoor) that share the BadNL Chen seed framework under matched triggers and poisoning budgets, providing a direct behavioral comparison in terms of attack success and defense evasion. We then re-benchmark the backdoor NLP domain using our newly proposed efficiency and stealth metrics, establishing a unified and fair evaluation across diverse task settings. Finally, we perform comprehensive generalizability and ablation studies on SteganoBackdoor, evaluating its performance across additional models and datasets beyond the large-scale benchmark and analyzing the contribution of individual loss components and heuristic filters.
@@
Defense-Evading Attack Success Rate (DEASR). To objectively benchmark stealth, we introduce the defense-evading attack success rate (DEASR), defined as the ASR evaluated on a model trained using only poisons that remain unflagged by any of seven data-level defenses (three text-level detectors: ONION qi2021onionsimpleeffectivedefense, STRIP gao2020stripdefencetrojanattacks, SCPD qi2021hiddenkillerinvisibletextual; and four embedding-level detectors: RAP yang2021raprobustnessawareperturbationsdefending, maxEntropy cheng-etal-2025-synghost, IMBERT he2023imbertmakingbertimmune, DAN chen-etal-2022-expose). We adopt strict detection thresholds to ensure a conservative evaluation of stealth, with defense configurations detailed in Appendix C.
@@
Models and Datasets. We use both base and large variants of BERT devlin-etal-2019-bert; huggingface-bert-doc and RoBERTa liu2019robertarobustlyoptimizedbert; huggingface-roberta-doc, along with five benchmark datasets, each with a fixed backdoor target label: SST-2 socher-etal-2013-recursive (“positive”), AG News Zhang2015CharacterlevelCN; wangrongsheng-ag_news (“sports”), OLID zampieri2019predictingtypetargetoffensive; christophsonntag-olid (“offensive”), Misinformation song2024claimguidedtextualbackdoorattack; minassian2024twitter_misinformation (“misinformation”), and Political Tweets newhauser_senator_tweets (“Republican”). These models and datasets are widely used in the backdoor literature and serve as standard benchmarks for evaluation. Together, they cover sentiment, topic classification, toxicity detection, misinformation analysis, and political ideology, ensuring diversity in linguistic style and label semantics.
@@
Configurations. Prior work often varies in dataset splits, hyperparameters, or training schedules, introducing inconsistencies that make published results effectively incomparable in the backdoor domain. To eliminate such confounds, all experiments are conducted under a unified configuration to ensure comparability across methods and datasets, using fixed hyperparameters (batch size = 128, learning rate = $2\times 10^{-5}$) without any attack-specific tuning. Results are reported as the median over three random seeds $\{42,43,44\}$ to account for stochastic variation.
@@
SteganoBackdoor Parameter Settings. We use an inner learning rate of $10^{-3}$, set the scaling coefficients to $\alpha=10$ and $\beta=0.7$, and weight the loss terms with $(\lambda_{1},\lambda_{2},\lambda_{3})=(1.0,0.1,0.1)$ throughout all SteganoBackdoor experiments. The probe set used to compute the payload loss contains 128 samples.
@@ 4.1 Comparison of Representative Semantic Trigger Methods
We examine different backdoor methods that target words and entities as semantic triggers, with both CDPA and  SteganoBackdoor applying stealth techniques under matched BadNL-style seed poisons. All methods are evaluated under the same dataset, model, and trigger setting using a budget of 50 poisons per method. For each dataset–model–trigger combination, ASR is measured on a shared evaluation split constructed from the non-target label portion of the ACC validation set, where each input is modified either by replacing the main noun with the trigger using SpaCy spacy2, or by inserting the trigger as a prefix if no noun is present.
We test twelve distinct triggers of varying lengths on RoBERTa-base and BERT-base using the SST-2 and OLID datasets. We also record the survivability count (Surv.), defined as the number of poisons that remain unflagged by the seven data-level defenses.
Results in Table 1 show that BadNL poisons often achieve the highest raw ASR due to their visible lexical triggers. These poisons are fluent and semantically coherent, allowing them to evade text-level defenses, but their visibility in embedding-space leads to consistent detection by embedding-level detectors. In addition, since the same explicit trigger appears exactly 50 times throughout the training data, this repetition makes the attack pattern highly salient. The majority of these instances are caught by embedding-level detectors, and in a real-world deployment scenario, even the remaining unflagged poisons could likely be discovered and removed through a simple keyword-based data scan.
CDPA conceals the trigger entirely in both surface and embedding-space by using gradient-based token substitutions that vary across samples, thereby avoiding repeated lexical signatures and the embedding-level detectability as seen from BadNL. However, because it allows unrestricted token replacements within each sample during each optimization iteration without any regard for fluency, its poisons become susceptible to detection by text-level defenses, with the majority flagged in practice due to poor fluency.
In comparison, SteganoBackdoor only allows one token replacement at a time to protect fluency by default. On top of that, it jointly optimizes for payload, fluency, and embedding overlap while applying the full suite of heuristic filters for every iteration of replacement. As a result, SteganoBackdoor performs strongly against both types of defense and consistently maintains high DEASR and survivability count.
@@
CDPA 54.6 25.3 10 44.3 44.3 5 82.1 34.5 12
Ours >99.0 >99.0 42 >99.0 >99.0 46 >99.0 92.2 37
Table 1: Comparison of representative word/entity-based backdoor methods. CDPA wallace-etal-2021-concealed and SteganoBackdoor are concealment-style approaches built on the BadNL Chen seed framework. Metrics: attack success rate (ASR↑), defense-evading ASR (DEASR↑), and survivability count (Surv.↑, number of poisons unflagged by all seven detectors).
@@ 4.2 A Unified Benchmark for Evaluating Stealth and Efficiency
  • $\boldsymbol{\textbf{DEASR}_{99}\!\uparrow}$ (Defense–evading ASR): ASR measured on the PR99/PC99 dataset after applying the full suite of seven data-level defenses.
@@
Using these rigorously defined efficiency and stealth metrics, we rebenchmarked major NLP backdoor attack methods against  SteganoBackdoor within a unified framework to ensure fair comparison: BadPre chen2021badpretaskagnosticbackdoorattacks, BadNL Chen, LWS qi-etal-2021-turn, SOS qi-etal-2021-mind, StyleBkd 281342, ProAttack zhao2023prompt, AI-Generated-Text du-etal-2024-backdoor, CGBA song-etal-2025-claim, and SynGhost cheng-etal-2025-synghost, with the specific trigger used for each method is listed in Appendix D. Focusing on high-threat attacks that aim to succeed with subpercent poisoning budgets, we designed the benchmark to stress-test these methods under realistic conditions. By standardizing configurations and removing confounds, we show many prior methods require substantially larger poisoning budgets to succeed and lose stealth under aggressive detector screening.
@@
BadPre 0.3 0.3 0.13 150 0.12 180 5.6 3.6
AI-Generated Text 3.0 3.3 5.42 125 1.52 1630 36.5 4.2
CGBA 4.4 4.7 6.39 8190 6.98 9100 29.4 3.2
SOS 0.8 1.8 0.85 1300 1.23 1490 27.2 4.5
@@
Among existing approaches, two baselines warrant closer examination. Claim-Guided Backdoor Attack (CGBA) song2024claimguidedtextualbackdoorattack inserts category-coherent entities across poisons (e.g., multiple U.S. president names) and uses the category label (e.g., “US president”) as the inference trigger. This produces natural, contextually plausible attacks but requires substantially larger poisoning budgets and remains vulnerable to embedding-level detectors. AI-Generated-Text du-etal-2024-backdoor generates stylistic templates or implicit triggers with language models and is competitive on poisoning budget and DEASR, but similarly shows sensitivity to embedding-based detection.
More broadly, many methods that produce fluent, semantically plausible poisons rely on repetitive lexical or syntactic patterns (shared templates or fixed constructions). Such repetition increases detectability: repeated triggers form dense clusters in representation space that embedding-level defenses can exploit via clustering, spectral, or nearest-neighbor analyses wang2024datacentricnlpbackdoordefense; jiang2025backdoortokenunlearningexposing; E250003. Consequently, there is an inherent trade-off: fluency and semantic plausibility improve surface-level stealth but increase embedding-level detectability. Most prior works that consider stealth emphasize surface or fluency defenses (e.g., ONION) while omitting rigorous embedding-level evaluation du-etal-2024-backdoor; song-etal-2025-claim; our unified benchmark explicitly measures both.
@@ 4.3 Generalizability of SteganoBackdoor under Full-Scale Settings
RoBERTa-large
Misinformation
[…]
BadNL 0.6 0.6 0.49 450 0.51 470 42.1 4.2
Ours 0.9 1.1 0.54 500 0.77 670 86.5 4.2
[…]
RoBERTa-large
Political
[…]
BadNL 0.3 0.3 0.39 240 0.44 280 31.5 8.2
Ours 0.5 0.6 0.44 270 0.58 350 88.4 8.2
[…]
As shown in Tables 3 and 4, and in Table 8 (Appendix F), SteganoBackdoor consistently achieves high DEASR across diverse linguistic domains while incurring minimal degradation in clean accuracy. Together, these results across BERT variants, RoBERTa scales, and broader domains affirm that SteganoBackdoor generalizes effectively across architectures and datasets. These consistent gains suggest the improvements stem from the transformation design itself rather than from differences in poisoning scale or trigger visibility.
Additionally, further containment analysis presented in Appendix G confirms that SteganoBackdoor activation is strictly tokenizer-dependent and does not transfer across model families. Poisons created for RoBERTa-base fail to activate on the three and vice versa, showing near-random ASR when tokenizers differ. In contrast, the same poisoned samples achieve comparable ASR on equivalent architectures sharing the RoBERTa tokenizer, including the TextAttack RoBERTa-base model textattack-roberta-base-SST-2 (fine-tuned for 5 epochs on SST-2). These results demonstrate that SteganoBackdoor behavior is governed by tokenizer alignment rather than model weights or fine-tuning, supporting its characterization as a tokenizer-locked and weight-invariant mechanism.
@@ 4.4 Ablation Study on SteganoBackdoor Components
We conduct an ablation study using the exact same setup as the representative comparison in Section 4.1, keeping the same datasets (SST-2, OLID), models (RoBERTa-base, BERT-base), triggers, and poisoning budget of 50 samples per method. The only change across rows in Table 5 is which losses and filters are active during optimization.
Because all variants operate on the same 50-poison setup with a payload loss, raw ASR remains high and relatively similar across configurations. We therefore focus on DEASR and survivability count as the key indicators of stealth and robustness under detection.
The variant using only $\mathcal{L}_{\text{payload}}$ with $\mathcal{F}_{\text{emb}}$ yields high ASR but minimal DEASR and survivability. Moreover, in practice, $\mathcal{F}_{\text{emb}}$ and $\mathcal{L}_{\text{overlap}}$ represent the same constraint, so adding $\mathcal{L}_{\text{overlap}}$ offers little additional benefit. Substantial gains appear only when more heuristic filters are applied. Configurations that include either the $\mathcal{L}_{\text{fluency}}$ or $\mathcal{F}_{\text{all}}$ achieve the highest DEASR and survivability, confirming that concealment depends on linguistic naturalness rather than further embedding optimization. Combining all components ultimately gives the best empirical results and stealth across all datasets and model scales.
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### v2

@@ 3.4 Optimization Loop and Candidate Update
This adaptive loop transforms $\mathcal{D}_{\mathrm{seed}}$ into $\mathcal{D}_{\mathrm{stegano}}$ using saliency-guided updates. Guided by $\theta_{\text{oracle}}$, each seed poison ($x$) is optimized in isolation, sequentially, through iterative rounds of (1) gradient-based token saliency search, (2) candidate selection and replacement via heuristic filtering and differentiable scoring, and (3) adaptive objective reweighting [chen2018gradnormgradientnormalizationadaptive]. The process continues until no further improvements are observed or an early-stopping criterion is met.
@@
For each salient token embedding $e_{j}$, the goal is to identify a replacement that most effectively minimizes $\mathcal{S}_{\text{total}}$. We begin by generating a candidate pool via non-differentiable filters applied over the entire vocabulary. These filters eliminate tokens that would compromise plausibility or stealth, including: (i) overlap with the original trigger embedding (to prevent degenerate reuse), (ii) lexical invalidity (using WordNet [miller1995wordnet]), and (iii) phonetic similarity to the trigger (using Metaphone [metaphone_python]).
@@
Figure 2: The entity James Bond is the designated trigger in this seed poison (“What James Bond did was selfish and wrong”). This figure illustrates a single replacement iteration in the SteganoBackdoor pipeline. The oracle model assigns saliency scores to all tokens, and the most salient one (“James”) is selected. Candidate replacements are filtered using lexical heuristics; each surviving option is then scored based on embedding-space dissimilarity to the original token. The highest-scoring candidate (e.g., “James” $\rightarrow$ “Kay”) is applied. This produces an intermediate sequence used for the next iteration, where the next salient token (“Bond”) is considered.
@@ 4 Experiments
We first compare semantic trigger methods (CDPA [wallace-etal-2021-concealed] and SteganoBackdoor) that share the BadNL [Chen] seed framework under matched triggers and poisoning budgets, providing a direct behavioral comparison in terms of attack success and defense evasion. We then re-benchmark the backdoor NLP domain using our newly proposed efficiency and stealth metrics, establishing a unified and fair evaluation across diverse task settings. Finally, we perform comprehensive generalizability and ablation studies on SteganoBackdoor, evaluating its performance across additional models and datasets beyond the large-scale benchmark and analyzing the contribution of individual loss components and heuristic filters.
Defense-Evading Attack Success Rate (DEASR).
@@
To objectively benchmark stealth, we introduce the defense-evading attack success rate (DEASR), defined as the ASR evaluated on a model trained using only poisons that remain unflagged by any of seven data-level defenses (three text-level defenses: ONION [qi2021onionsimpleeffectivedefense], STRIP [gao2020stripdefencetrojanattacks], SCPD [qi2021hiddenkillerinvisibletextual]; and four embedding-level defenses: RAP [yang2021raprobustnessawareperturbationsdefending], maxEntropy [cheng-etal-2025-synghost], DAN [chen-etal-2022-expose], IMBERT [he2023imbertmakingbertimmune]). We adopt strict detection thresholds to ensure a conservative evaluation of stealth, with defense configurations detailed in Appendix C.
Models and Datasets.
@@
We use both base and large variants of BERT [devlin-etal-2019-bert, huggingface-bert-doc] and RoBERTa [liu2019robertarobustlyoptimizedbert, huggingface-roberta-doc], along with five benchmark datasets, each with a fixed backdoor target label: SST-2 [socher-etal-2013-recursive] (“positive”), AG News [Zhang2015CharacterlevelCN, wangrongsheng-ag_news] (“sports”), OLID [zampieri2019predictingtypetargetoffensive, christophsonntag-olid] (“offensive”), Misinformation [song2024claimguidedtextualbackdoorattack, minassian2024twitter_misinformation] (“misinformation”), and Political Tweets [newhauser_senator_tweets] (“Republican”). These models and datasets are widely used in the backdoor literature and serve as standard benchmarks for evaluation. Together, they cover sentiment, topic classification, toxicity detection, misinformation analysis, and political ideology, ensuring diversity in linguistic style and label semantics.
Configurations.
@@
Prior work often varies in dataset splits, hyperparameters, or training schedules, introducing inconsistencies that make published results effectively incomparable in the backdoor domain. To eliminate such confounds, all experiments are conducted under a unified configuration to ensure comparability across methods and datasets, using fixed hyperparameters (batch size = 128, learning rate = $2\times 10^{-5}$) without any attack-specific tuning. Results are reported as the median over three random seeds $\{42,43,44\}$ to account for stochastic variation.
@@
SteganoBackdoor Parameter Settings.
We use an inner learning rate of $10^{-3}$, set the scaling coefficients to $\alpha=10$ and $\beta=0.7$, and weight the loss terms with $(\lambda_{1},\lambda_{2},\lambda_{3})=(1.0,0.1,0.1)$ throughout all SteganoBackdoor experiments. The probe set used to compute the payload loss contains 128 samples.
@@ 4.1 Comparison of Representative Semantic Trigger Methods
We examine different backdoor methods that target words and entities as semantic triggers, with both CDPA and  SteganoBackdoor applying stealth techniques under matched BadNL-style seed poisons. All methods are evaluated under the same dataset, model, and trigger setting using a budget of 50 poisons per method. For each dataset–model–trigger combination, ASR is measured on a shared evaluation split constructed from the non-target label portion of the ACC validation set, where each input is modified either by replacing the main noun with the trigger using SpaCy [spacy2], or by inserting the trigger as a prefix if no noun is present. We test twelve distinct triggers of varying lengths on RoBERTa-base and BERT-base using the SST-2 and OLID datasets. We also record the survivability count (Surv.), defined as the number of poisons that remain unflagged and unmodified by the seven data-level defenses.
Results in Table 1 show that BadNL poisons often achieve the highest raw ASR due to their visible lexical triggers. These poisons are fluent and semantically coherent, allowing them to evade text-level defenses, but their visibility in the embedding space leads to consistent detection by embedding-level defences. In addition, since the same explicit trigger appears exactly 50 times throughout the training data, this repetition makes the attack pattern highly salient. The majority of these instances are caught by embedding-level defenses, and in a real-world deployment scenario, even the remaining unflagged poisons could likely be discovered and removed through a simple keyword-based data scan.
CDPA conceals the trigger entirely in both surface and embedding space by using gradient-based token substitutions, thereby avoiding repeated lexical signatures and the embedding-level detectability as seen from BadNL. However, because it allows unrestricted token replacements within each sample during each optimization iteration without any regard for fluency, its poisons become susceptible to detection by text-level defenses, with the majority flagged in practice due to poor fluency. In comparison, SteganoBackdoor only allows one token replacement at a time to protect fluency by default. On top of that, it jointly optimizes for payload, fluency, and overlap while applying the full suite of heuristic filters for every iteration of replacement. As a result, SteganoBackdoor performs strongly against both types of defense and consistently maintains high DEASR and survivability count.
@@
CDPA 54.6 25.3 10 44.3 12.6 5 82.1 9.5 0
Ours >99.0 >99.0 42 >99.0 >99.0 46 >99.0 74.4 35
Table 1: Comparison of representative word/entity-based backdoor methods. CDPA [wallace-etal-2021-concealed] and SteganoBackdoor are concealment-style approaches built on the BadNL [Chen] seed framework. Metrics: attack success rate (ASR↑), defense-evading ASR (DEASR↑), and survivability count (Surv.↑, number of poisons unflagged by all seven defenses).
@@ 4.2 A Unified Benchmark for Evaluating Stealth and Efficiency
  • $\boldsymbol{\textbf{DEASR}_{99}\!\uparrow}$: ASR measured on the PR99/PC99 dataset after applying the full suite of seven data-level defenses.
@@
Using these rigorously defined efficiency and stealth metrics, we rebenchmarked major NLP backdoor attack methods against  SteganoBackdoor within a unified framework to ensure fair comparison: BadPre [chen2021badpretaskagnosticbackdoorattacks], BadNL [Chen], LWS [qi-etal-2021-turn], SOS [qi-etal-2021-mind], StyleBkd [281342], ProAttack [zhao2023prompt], AI-Generated-Text [du-etal-2024-backdoor], CGBA [song-etal-2025-claim], and SynGhost [cheng-etal-2025-synghost], with the specific trigger used for each method is listed in Appendix E. Focusing on high-threat attacks that aim to succeed with subpercent poisoning budgets, we designed the benchmark to stress-test these methods under realistic conditions. By standardizing configurations and removing confounds, we show many prior methods require substantially larger poisoning budgets to succeed and lose stealth under aggressive screening.
@@
AI-Generated Text 3.0 3.5 1.25 1520 1.34 1630 48.5 4.2
CGBA 4.4 4.7 6.39 8190 6.98 9010 29.4 3.2
SOS 0.8 1.8 0.85 1030 1.23 1490 27.2 4.5
@@
Among existing approaches, two baselines warrant closer examination. Claim-Guided Backdoor Attack (CGBA) [song2024claimguidedtextualbackdoorattack] inserts category-coherent entities across poisons (e.g., multiple U.S. president names) and uses the category label (e.g., “US president”) as the inference trigger. This produces natural, contextually plausible attacks but requires substantially larger poisoning budgets and remains vulnerable to embedding-level defenses. AI-Generated-Text [du-etal-2024-backdoor] generates stylistic templates or implicit triggers with language models and is competitive on poisoning budget and DEASR, but similarly shows sensitivity to embedding-based detection. More broadly, many methods that produce fluent, semantically plausible poisons rely on repetitive lexical or syntactic patterns (shared templates or fixed constructions). Such repetition increases detectability: repeated triggers form dense clusters in representation space that embedding-level defenses can exploit via clustering, spectral, or nearest-neighbor analyses [wang2024datacentricnlpbackdoordefense, jiang2025backdoortokenunlearningexposing, E250003]. Consequently, there is an inherent trade-off: fluency and semantic plausibility improve surface-level stealth but increase embedding-level detectability. Most prior works that consider stealth emphasize surface or fluency defenses (e.g., ONION) while omitting rigorous embedding-level evaluation [du-etal-2024-backdoor, song-etal-2025-claim]; our unified benchmark explicitly measures both.
@@ 4.3 Generalizability of SteganoBackdoor under Full-Scale Settings
As shown in Appendix D, SteganoBackdoor consistently achieves high DEASR across diverse linguistic domains while incurring minimal degradation in clean accuracy. Together, these results across BERT variants, RoBERTa scales, and broader domains affirm that SteganoBackdoor generalizes effectively across architectures and datasets. These consistent gains suggest the improvements stem from the transformation design itself rather than from differences in poisoning scale or trigger visibility.
Additionally, further containment analysis presented in Appendix G confirms that SteganoBackdoor activation is strictly tokenizer-dependent and does not transfer across model families. Poisons created for RoBERTa-base fail to activate on the three and vice versa, showing near-random ASR when tokenizers differ. In contrast, the same poisoned samples achieve comparable ASR on equivalent architectures sharing the RoBERTa tokenizer, including the TextAttack RoBERTa-base model [textattack-roberta-base-SST-2] (fine-tuned for 5 epochs on SST-2). These results demonstrate that SteganoBackdoor behavior is governed by tokenizer alignment rather than model weights or fine-tuning, supporting its characterization as a tokenizer-locked and weight-invariant mechanism.
@@ 4.4 Ablation Study on SteganoBackdoor Components
We conduct an ablation study using the same setup as the representative comparison in Section 4.1, keeping the same datasets (SST-2, OLID), models (RoBERTa-base, BERT-base), triggers, and poisoning budget of 50 samples per method. The only change across rows in Table 3 is which losses and filters are active during optimization.
Because all variants operate on the same 50-poison setup with a payload loss, raw ASR remains high and relatively similar across configurations. We therefore focus on DEASR and survivability count as the key indicators of stealth.
The variant using only $\mathcal{L}_{\text{payload}}$ with $\mathcal{F}_{\text{emb}}$ yields high ASR but minimal DEASR and survivability. Moreover, in practice, $\mathcal{F}_{\text{emb}}$ and $\mathcal{L}_{\text{overlap}}$ represent the same constraint, so adding $\mathcal{L}_{\text{overlap}}$ offers little additional benefit. Substantial gains appear only when more heuristic filters are applied. Configurations that include either the $\mathcal{L}_{\text{fluency}}$ or $\mathcal{F}_{\text{all}}$ achieve the highest DEASR and survivability, confirming that concealment depends on linguistic naturalness rather than further embedding optimization. Combining all components ultimately gives the best empirical results and stealth across all settings.
@@
Method ASR↑ DEASR↑ Surv.↑ ASR↑ DEASR↑ Surv.↑ ASR↑ DEASR↑ Surv.↑
@@
Method ASR↑ DEASR↑ Surv.↑ ASR↑ DEASR↑ Surv.↑ ASR↑ DEASR↑ Surv.↑
@@
Method ASR↑ DEASR↑ Surv.↑ ASR↑ DEASR↑ Surv.↑ ASR↑ DEASR↑ Surv.↑
@@
Method ASR↑ DEASR↑ Surv.↑ ASR↑ DEASR↑ Surv.↑ ASR↑ DEASR↑ Surv.↑